3. For a long time as a younger adult, a certain physics instructors eyes each had the following focal range: FP- 35.0 cm NP 12.0 cm a. What was the Refractive Power (in Diopters) of the glasses he wore to allow him standard far viewing? b. How close could he hold a book to his eyes (while wearing these glasses) and still focus on the words in the book? Then time passed. Alas. Now his focal range is this: FP 25.0 cm NP 17.0 cm What is the Refractive Power (in Diopters) of the glasses he now needs so that he still has standard far viewing d. How close can he hold a book to his eyes (while wearing these glasses) and still focus on the words in the book? (After you complete result d, notice that its greater than the standard near-viewing distance. This is why he had to switch to bi-focals-two different sets of lenses: one for distance viewing, the other for close viewing.)
Answer #1

object distance s = infinity


image distance s' = -25 cm


focal length = f


1/s + 1/s' = 1/f

f = -25 cm

power = 1/f (in m) = -1/0.25 = -4 D

---------------


(d)


for near viewing

object dsistance s = ?

image distance s' = -17 cm

focal length f = -25 cm


1/s - 1/17 = -1/25


s = 53.1 cm <<<------------answer